A tuning fork of frequency of 500 c/s is sounded on resonance tube. the first and second resonance are obtained at 17 cm and 50 cm. the velocity of sound in m/sec is
(a) 170
(b) 350
(c) 520
(d) 850

Grade:Upto college level

2 Answers

Arun
3 years ago

The first resonance occurs at, while the second resonance occurs at.

Therefore,

and.

If we subtract the two equations from each other:

.

Hence,

 

optionB is correct

Vikas TU
3 years ago
The first resonance occurs at lambda/4, while the second resonance occurs at 3lambda/4.
Therefore,
17+e=lambda/4 and 52+e=3lambda/4.
If we subtract the two equations from each other:
52-17= lambda/
lambda=70cm = 0.7m.
Hence,
v=f*lambda
v=500*0.7=350 m/s